OpenStudy (anonymous):

Evaluate the expression 5(7 − 4)^2 ÷ 3 + 11. 3 14 26 33

OpenStudy (briellshaw):

you have to use PEMDAS

OpenStudy (briellshaw):

Parentheses exponents multiplication division addition subtraction

OpenStudy (anonymous):

5(3)2/3+11

OpenStudy (briellshaw):

21

OpenStudy (briellshaw):

5*3=15*2=3=/3=10+11=21

zepdrix (zepdrix):

\[\rm 5(7-4)^2\div3+11\]\[\rm 5(3)^2\div3+11\]Your first step looks good :)

OpenStudy (anonymous):

the 2 is an exponet

OpenStudy (anonymous):

\[5\times9\div3+11\]

OpenStudy (texaschic101):

now do multiplication

OpenStudy (texaschic101):

5 * 9 = ?